A tennis ball is thrown vertically upward with an initial velocity of +8.0cm/s.
a.What will the balls speed be when it returns to its starting point.
b.How long will the ball take to reach its starting point?

Answer by Alan3354(69443)

A tennis ball is thrown vertically upward with an initial velocity of +8.0cm/s.
a.What will the balls speed be when it returns to its starting point.
It's the same, 8 cm/sec
===========================
b.How long will the ball take to reach its starting point?
--------
Using 10 m/sec/sec for gravity:
h(t) = -5t^2 + 0.08t
Start point is at h(t) = 0
-5t^2 + 0.08t = 0
t*(-5t + 0.08) = 0
t = 0 -- at launch
t = 0.08/5 = 0.016 seconds